A cannon ball is fired with an initial velocity of 25m/s at 53º from the horizontal. The cannon is
situated on a cliff 60 m high. Find the (a) maximum height the cannon ball reached; (b) the
maximum horizontal distance it reached; (c) total time the cannon ball is in the air.
1
Expert's answer
2020-11-16T07:48:06-0500

a) Let's write the equations of motion of the cannon ball in horizontal and vertical directions:


"x=v_0tcos\theta, (1)""y=y_0+v_0tsin\theta-\dfrac{1}{2}gt^2, (2)"

here, "x" is the horizontal displacement of the cannon ball (or the range of the cannon ball), "v_0=25.0\ \dfrac{m}{s}" is the initial velocity of the cannon ball, "t" is the total flight time of the cannon ball, "\theta=53^{\circ}" is the launch angle, "y" is the vertical displacement of the cannon ball (or the height), "y_0=60.0\ m" is the initial height from which the cannon ball was fired and "g=9.8\ \dfrac{m}{s^2}" is the acceleration due to gravity.

Let's first find the time that the cannon ball takes to reach the maximum height from the kinematic equation:


"v_y=v_0sin\theta-gt_{rise},""0=v_0sin\theta-gt_{rise},""t_{rise}=\dfrac{v_0sin\theta}{g}."

Then, we can substitute "t_{rise}" into the second equation and find the maximum height:


"y_{max}=y_0+v_0sin\theta\cdot\dfrac{v_0sin\theta}{g}-\dfrac{1}{2}g(\dfrac{v_0sin\theta}{g})^2,""y_{max}=y_0+\dfrac{v_0^2sin^2\theta}{2g},""y_{max}=60\ m+\dfrac{(25.0\ \dfrac{m}{s})^2\cdot sin^253^{\circ}}{2\cdot 9.8\ \dfrac{m}{s^2}}=80.3\ m."


(b)-(c) Let's first find total flight time of the cannon ball:


"y=y_0+v_0tsin\theta-\dfrac{1}{2}gt^2,""0=60+25tsin53^{\circ}-\dfrac{1}{2}\cdot 9.8t^2,""4.9t2-20t-60=0."

This quadratic equation has two roots:


"t_1=\dfrac{-b+\sqrt{b^2-4ac}}{2a}=\dfrac{20+\sqrt{(-20)^2+1176}}{2\cdot 4.9}=6.09\ s,""t_2=\dfrac{-b-\sqrt{b^2-4ac}}{2a}=\dfrac{20-\sqrt{(-20)^2+1176}}{2\cdot 4.9}=-2.01\ s."


Since time can't be negative, the correct answer is "t=6.09\ s."

Finall, we can substitute "t" into the first equation and find the maximum horizontal distance reached by the cannon ball:


"x=v_0tcos\theta=25.0\ \dfrac{m}{s}\cdot 6.09\ s\cdot cos53^{\circ}=91.6\ m."

Answer:

(a) "y_{max}=80.3\ m."

(b) "x=91.6\ m."

(c) "t=6.09\ s."